OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12)

OpenStudy (anonymous):

Find the slope first... Slope(m) = (12-7)/(3+2) = 5/5 = 1 m = 1 Now the required equation is: \[y - y_1 = m(x - x_1)\] y - 12 = 1(x - 3) y = x + 9